OpenStudy (anonymous):

One day HINDU and MUSLIM were on their way. They saw that ox was smelling dung. Now, seeing that, MUSLIM told to HINDU that his father (according to hindu culture) was eating dung. After listening such, Hindu felt same as we felt as we are hindus. But, using correct time and situation, HINDU told something to MUSLIM that he kept quiet and can't say a word to HINDU. Now, the question is - just think and tell what might had the HINDU said to MUSLIM....!!

OpenStudy (anonymous):

theyre both stupid

OpenStudy (anonymous):

religion is dumb, where is the evidence for god or miracles?

OpenStudy (anonymous):

sathi, whats about your sequence or series

OpenStudy (anonymous):

3, 7, 32, 211, ..

OpenStudy (anonymous):

nobody know ;(

OpenStudy (anonymous):

what about religion, why you believe in fairy tales?

OpenStudy (anonymous):

Cantor that's a little harsh. But so, so, true (and hilarious).

OpenStudy (anonymous):

nobody could solve the numerical problems so am tired........of asking such numerical and just to entertain

OpenStudy (anonymous):

newton, im stumped here with this diffq q

OpenStudy (anonymous):

the website has this answer http://www.sosmath.com/diffeq/first/application/population/population.html

OpenStudy (anonymous):

P(t) = MCe^(kt)/ ( M + Ce^kt) i got P = P(t) = MCe^(kt)/ ( 1 + Ce^kt)

OpenStudy (anonymous):

website has P(t) = MCe^(kt)/ ( M + Ce^kt)

OpenStudy (anonymous):

i outlined my solution in the previous post. both ways seem logical, but oddly enough we get two different solutions

OpenStudy (anonymous):

well m not the student of math but preparing for medicine and for medicine these questions are asked in the exam

OpenStudy (anonymous):

Where is your working?

OpenStudy (anonymous):

previous post

OpenStudy (anonymous):

one down from the hindu cattle problem

OpenStudy (anonymous):

no...not that

OpenStudy (anonymous):

sathi post your sequence problem again, i will solve it

OpenStudy (anonymous):

Ah, sorry, I see

OpenStudy (anonymous):

3,7,32,211, ?, ?

OpenStudy (anonymous):

ok one sec

OpenStudy (anonymous):

i guess u have already taken more than 1 sec lol

OpenStudy (anonymous):

im looking for calculator

OpenStudy (anonymous):

theres no pattern

OpenStudy (anonymous):

3,7 and 211 are prime numbers

OpenStudy (anonymous):

i see 7 = 3 + 2^2, where 2 is the prime less than 3

OpenStudy (anonymous):

32 = 7 + 5^2 , where 5 is one prime less than 7

OpenStudy (anonymous):

whats the answer?

OpenStudy (anonymous):

you got the sequence wrong, its 3, 7, 31, 211